A series RLC circuit has R = 343Ω, L = 1.8H, C = 5µF. It is connected to an AC source with f = 60.0 Hz and ΔVmax = 150 V. What is the maximum current through the circuit? Please show work.
Answer #1

net impedance of the circuit

Z^2 = R^2 + ( 2 pi f L - 1/ ( 2 pi fC)) ^2

Z = 373.352 ohm

max current through the circuit

i = Vo/ Z = 150 / 373.352

i = 0.4018 A

==========

Comment before rate in case any doubt, will reply for sure.. goodluck